# Reviewing Scientific Papers for Critical Problems With Reasoning LLMs: Baseline Approaches and Automatic Evaluation

### v1

@@ 5 Limitation
As a proof-of-concept study, this work has several limitations. First, only closed-source reasoning LLMs were evaluated. Future work can consider comparing different PDF preprocessing pipelines and open-source LLMs. Second, our evaluation metrics may suffer inaccuracies and inflation due to (1) ambiguities in paper authors’ retraction comments (e.g., "withdrawn due to a crucial error in Lemma 2", but what exactly is the error?), (2) automatic evaluation utilizing LLMs without human involvement, and (3) our settings that each LLM quality checker was tested only once per paper and each LLM judge graded each submission only once. Nevertheless, a higher score under parallel evaluation still reasonably indicates better performance. We expect gradual improvements in the reliability of the LLMs-as-judges approach as leading models continue to evolve. Third, the influence of experiment parameters or alternate prompts were not investigated. In addition, our results based on a dataset rich in math and physics papers published in the past may not generalize well to papers in other scientific domains or future papers. We plan to address this limitation in future work. Last, since we were unable to recruit domain experts for manual annotation, it remains possible that some cases in the dataset contain problems that could not be detected based on the manuscript alone, or that the authors gave inaccurate reasons for retraction, which may pose an upper limit for model performance.
We noticed a concurrent work uploaded to arXiv on May 17 by Son et al. (2025), which also utilized the WithdrarXiv dataset for LLM-driven automatic error detection in scientific papers. Unlike their work which focused mainly on benchmarking LLMs, our work approached the same topic more from an application perspective. Their main metric overlaps substantially with our hit rate, although there is a difference in the meaning of $k$. Other key differences include: (1) Son et al. present a small but better annotated dataset designed solely for benchmarking, whereas our dataset is much larger, including a training set which may directly benefit future work; (2) they normalized the papers using OCR, whereas we assessed LLMs with PDF and LaTeX; (3) they used GPT-4.1 to determine whether an LLM’s error submissions match the gold error annotation, whereas we adopted multiple reasoning LLMs as judges. The strengths of their study include the use of domain experts for further quality control of their benchmark, and the use of resampling to study variations in LLM outputs. The two efforts are complementary, and both report the leading performance of o3 on scientific verification.
@@
We will publish the WithdrarXiv-Check dataset, model outputs (including thinking outputs if available), and core experiment code to support future work by the research community. We will update this paper with the link to our public repository in the near future.
@@ Acknowledgments
We thank community members of the Black Spatula Project for their inspiring discussions in December 2024, especially Ugo Dos Reis for highlighting the WithdrarXiv dataset, and Delip Rao for instructions on how to use this dataset. The authors also thank Yuxin Xu for her help with the API.
We welcome discussions, collaborations, and funding to further improve this work.
@@ Appendix A Prompt Templates
Did my colleague find exactly the same problem that the authors mentioned in the retraction comment? In your final answer, simply give "Yes" or "No". Make your best decision if you are unsure. You may explain your decision but please be concise.
Checker Model Judge Model
Gemini 2.5 Pro o3 (med)
PDF as an attachment
Gemini 2.5 Pro 64.9 62.4
Gemini 2.5 Flash 62.0 64.1
o3 (medium) 72.7 75.5
o4-mini (medium) 69.0 68.2
Claude 3.7 Sonnet 21.6 19.2
LaTeX script in prompt
Gemini 2.5 Pro 58.8 65.3
Gemini 2.5 Flash 58.8 61.6
o3 (medium) 75.5 80.4
o4-mini (medium) 66.1 71.4
Claude 3.7 Sonnet 44.5 40.8
Table 3: Hit rate results by individual judge.
@@ Appendix B Model Parameters
tools: []
temperature: 0
seed: 42
@@
reasoning summary: "detailed"

### v2

@@ 5 Limitation
As a proof-of-concept study, this work has several limitations. First, only closed-source reasoning LLMs were evaluated. Future work may consider comparing different PDF preprocessing pipelines and open-source LLMs. Second, our evaluation metrics may suffer inaccuracies due to (1) ambiguities in paper authors’ retraction comments (e.g., "withdrawn due to a crucial error in Lemma 2", without specification of the error, (2) automatic evaluation utilizing LLMs without human involvement, and (3) our settings that each LLM checker was tested only once per paper and each LLM judge graded each submission only once. Nevertheless, a higher score under parallel evaluation still reasonably indicates better performance of an LLM checker. We expect gradual improvements in the reliability of the LLMs-as-judges approach as leading models continue to evolve. Third, the impacts of experiment parameters or alternate prompts were not formally investigated or optimized. In addition, our results based on a dataset rich in math and physics papers published in the past may not generalize well to papers in other scientific domains or future papers. Last, since we were unable to recruit domain experts for manual annotation, it remains possible that some cases in the dataset contain problems that could not be detected based on the manuscript alone, or that the authors gave inaccurate reasons for retraction, which may pose an upper limit for model performance. If conditions permit, future work should also consider involving domain experts in calibrating the LLM judges to resolve potential biases caused by circular evaluation of LLMs using LLMs.
A concurrent work on arXiv by Son et al. (2025) also utilizes the WithdrarXiv dataset for automatic error detection in scientific papers. Unlike their work which focuses on benchmarking LLMs, our work approaches the same topic more from an application perspective. Their main metrics overlap substantially with our hit rate, although there is a difference in the meaning of $k$. Other key differences include:
  • Son et al. present a small but better annotated dataset of latest papers designed solely for one-time benchmarking, whereas our dataset is much larger and contains more papers from the past, also including a training set which may directly benefit future work.
  • They normalized all papers using OCR, whereas we assessed LLMs with papers in PDF and LaTeX formats.
  • They used GPT-4.1 to determine whether an LLM’s error submissions match the gold error annotation, whereas we adopted multiple reasoning LLMs as judges.
  • They treated all LLM-identified problems that did not directly match the gold error description as wrong answers, whereas we allowed some flexibility through LLM reasoning and attempted to assess the precision of these additional answers.
These factors could explain the gap between their and our evaluation results. The strengths of their study include the use of domain experts for further quality control of their benchmark dataset, inclusion of open-source LLMs in evaluation, and a study of variations in LLM outputs via resampling. The two efforts are complementary, and both report the leading performance of o3.
@@
Our WithdrarXiv-Check dataset, experiment code, and model outputs (including thinking outputs if available) are available on Github [^10] https://github.com/tianmai-zhang/WithdrarXiv-Check. Readers may use the dataset for further experiments, reproduce the numbers in our tables, or perform a closer inspection of the model outputs to obtain further insights into LLM behaviors. We welcome discussions, collaborations, and funding for future work.
@@ Acknowledgments
We thank community members of the Black Spatula Project for their inspiring discussions, especially Ugo Dos Reis for highlighting the WithdrarXiv dataset in December 2024, Delip Rao for instructions on how to use this dataset, and Casey Wimsatt and David Reinstein for constructive feedback on early versions of this paper. The authors also thank Yuxin Xu for her help with the API.
@@ Appendix A Prompt Templates
Is my colleague referring to exactly the same problem mentioned in the retraction comment? Your final answer should be "Yes" or "No". Default your answer to "No" and only give "Yes" if you are certain. You may explain your decision but please be concise.
Prompt for LLM judges to determine true positives:
My colleague was reading this paper and said there is a critical problem in it, as described below:
Problem: {problem}
Location: {location}
Explanation: {explanation}
Is this problem a true problem or a false alarm? Please be careful because I don’t want to get the authors into trouble by mistake. In your final answer, clearly indicate "Yes, it is a true problem" or "No, it is a false alarm". Make your best decision if you are unsure. You may explain your decision but please be concise.
{paper_attachment}
@@ Appendix B Model Parameters
include thoughts: True
tools: []  temperature: 0  seed: 42
@@
reasoning summary: "auto"